OpenStudy (anonymous):

Which of the following equations respresents a line parallel to y = 5x - 7? Question 6 options: 9y = 5x + 2 9y -45x = -7 y = 45x - 63 9y = 5x - 7

OpenStudy (anonymous):

y = 45x - 63

OpenStudy (aroub):

9y -45x = -7

OpenStudy (aroub):

parallel=same slope

OpenStudy (anonymous):

I came up with the same as aroub.

OpenStudy (anonymous):

a line should be in the form of y=mx+b that answer is not right because if you solv for y the slope will be different

OpenStudy (aroub):

It does not matter.

OpenStudy (aroub):

What matters is the slope, the slope has to be the same.

OpenStudy (anonymous):

if it is not in the form of y=mx+b you cannot call it the slope it is a different equation

OpenStudy (anonymous):

right the slope does has to be the same

OpenStudy (aroub):

9y-45x=-7 return it to the form of y=mx+b => 9y=45x-7 y=5x-7/9 ^ see! The slope is the same :)

OpenStudy (anonymous):

I knew that was the answer.

OpenStudy (anonymous):

your are right my mistake

OpenStudy (aroub):

No problem!

OpenStudy (anonymous):

So problem solved!

OpenStudy (aroub):

^ Yup :D